Could somebody please help me with scientific notation? I don't understand what you are supposed to do and the steps in which you do it. My teacher isn't very good at explaining things (he's a grouch :)) and I don't want to hurt my grade because I don't get it. I don't exactly have a specific problem that I need worked out, so example problems are just fine too. Thanks so much,
A really appreciative 9th grader. . .

Found 3 solutions by Alan3354, richwmiller, stanbon:
Answer by Alan3354(69443)

google it. 929,000 entries there

Answer by richwmiller(17219)

Assuming that you are reporting the 100% truth, which I doubt. The teacher may well be incompetent.
either you or your parents should talk with a vice- principal or ombudsman or counselor about this "grouch". Since you are in 9th grade, there should be a vice principal or counselor to talk to.
That is unacceptable.
So give an example of your problem with scientific notation.
6,234=6.234*10^3
6000=6*10^3
1000=10^3
How many zeroes? 3 what power 3

Answer by stanbon(75887)

2.35x10^4 means "move the decimal point 4 to the right" to get 23,500
---
2.35x10^-4 means "move the decimal point 4 to the left" to get 0.000235
--------------------------
========================
The scientific notation form
for 6045199 = 6.045199x10^6 ; the 6 tells you you would have to move
the decimal point 6 to the right to get the original number.
---
for 0.0001468 = 1.468x10^-4 ; the -4 tells you you would have to move
the decimal point 4 to the left to get the original number.
===============================================================
Notice that each of the scientific examples is expressed as a
number greater than one and less than 10 followed by some power
of ten.
-------------------
Google will tell you more about scientific notation if you want
to read it.
-----------------------------
Cheers,
Stan H.